Wrte the numbers that point j and k represent on the number line
20 J K 110
This is a number line I have to fill in the missing numbers

Answer by CubeyThePenguin(3113)

On a number line, ticks are evenly spaced.

interval = (110 - 20)/3 = 90/3 = 30

20, 50, 80, 110

J = 50
K = 80